Consider the equation xe^x = cos x


(a) Apply the intermediate value theorem to show that the function has a root in the interval


[0, 1].




1
Expert's answer
2022-05-18T09:10:20-0400

Function "f(x)=xe^x-\cos x" is continuous on "[0, 1]."


"f(0)=0(e^0)-\cos(0)=-1<0"

"f(1)=1(e^1)-\cos(1)=e-\cos (1)>0,"

since "e>1, \cos(1)<1."

Then by the intermediate value theorem there exists the number "c" in "(0, 1)" such that "f(c)=0."

Therefore the function "f(x)=xe^x-\cos x" has a root in the interval "[0, 1]."